# DiffGAD: A Diffusion-based Unsupervised Graph Anomaly Detector

### v1

@@ C.1 Detailed Dataset Descriptions
Reddit (Kumar et al., 2019; Wang et al., 2021): Reddit consists of the user posts on subreddits within one month, and the text of each post is transferred to a feature vector to represent the LIWC categories (Pennebaker et al., 2001). The features of users and subreddits are derived by summing the features of their respective posts, where the banned users on the platform are considered to be anomalous.
@@
C.2 More Experimental Comparisons
In this subsection, we show more experimental metrics, including Average Precision (AP), Recall@k, and Area Under the Precision and Recall Curve (AUPRC), and the experimental results are listed in Table 6, Table 7, and Table 8, respectively. Specifically, we can observe that our DiffGAD achieves favorable performances with small $std$ results around all benchmarks, which demonstrates our effectiveness.
@@ C.2 More Experimental Comparisons
C.3 More Ablation Results
In this subsection, we provide more ablation results to validate our method. Specifically, we empirically explore the validity of the generated common feature $c_{next}$, and show the distance between $c_{next}$ with the reconstructed Abnormal and Normal features statistically, the L2 distances is shown in Table 9, and the Standardized cosine distances in depicted in Table 10. We can observe that $c_{next}$ is more similar to the normal samples from both results, which shows that normal samples contain more common attributes.
@@ C.3 More Ablation Results
C.4 Detailed Baseline Description
@@
DOMINANT (Ding et al., 2019): DOMINANT is one of the pioneering works that leverages GCN and AE for anomaly detection. It follows an encoder-decoder architecture with a two-layer GCN. The node feature decoder is also a two-layer GCN, while the graph structure is decoded by a one-layer GCN and dot product. The anomaly score for each node is determined by the weighted sum of both decoders.

### v2

@@ C.2 Detailed Dataset Descriptions
Reddit (Kumar et al., 2019; Wang et al., 2021): Reddit consists of user posts on subreddits within one month, and the text of each post is transferred to a feature vector to represent the LIWC categories (Pennebaker et al., 2001). The features of users and subreddits are derived by summing the features of respective posts, where the banned users on the platform are considered to be anomalous.
@@
C.3 More Experimental Comparisons
In this subsection, we show more experimental metrics, including Average Precision (AP), Recall@k, and Area Under the Precision and Recall Curve (AUPRC), and the experimental results are listed in Table 6,  7, and 8, respectively. Specifically, we can observe that our DiffGAD achieves favorable performances with small $std$ results around all benchmarks, which demonstrates our effectiveness.
@@ C.3 More Experimental Comparisons
C.4 More Ablation Results
In this subsection, we provide more ablation results to validate our method. Specifically, we empirically explore the validity of the generated common feature $c_{next}$, and show the distance between $c_{next}$ with the reconstructed Abnormal and Normal features statistically, the L2 distances are shown in Table 9, and the Standardized cosine distances are depicted in Table 10. We can observe that $c_{next}$ is more similar to the normal samples from both results, which shows that normal samples contain more common attributes.
@@ C.4 More Ablation Results
C.5 Adaptation with different backbones.
Table 11: ROC-AUC result of different encoder-encoder architectures on 5 datasets, where we show the avg perf. $\pm$ the std of perf. of each method. The best results of all methods are indicated in boldface.
Backbone Weibo Reddit Disney Books Enron Avg.
MLP 85.5 $\pm$ 0.3 50.6 $\pm$ 0.0 48.1 $\pm$ 0.1 49.3 $\pm$ 5.9 41.3 $\pm$ 14.1 55.0 $\pm$ 4.1
VAE 84.4 $\pm$ 0.2 50.6 $\pm$ 0.0 48.0 $\pm$ 0.1 50.5 $\pm$ 5.4 42.9 $\pm$ 8.2 55.3 $\pm$ 2.8
VGAE 92.5 $\pm$ 0.0 55.8 $\pm$ 0.1 47.8 $\pm$ 3.3 48.4 $\pm$ 4.2 56.8 $\pm$ 0.7 60.3 $\pm$ 1.7
GTrans 90.9 $\pm$ 0.4 56.1 $\pm$ 0.0 31.0 $\pm$ 7.0 32.5 $\pm$ 12.3 55.5 $\pm$ 0.6 53.2 $\pm$ 4.1
GAE 92.0 $\pm$ 0.1 56.1 $\pm$ 0.0 40.3 $\pm$ 6.7 59.1 $\pm$ 2.5 59.1 $\pm$ 1.6 61.3 $\pm$ 2.2
Table 12: ROC-AUC result of DiffGAD with different encoder-encoder architectures on 5 datasets, where we show the avg perf. $\pm$ the std of perf. of each method. The best results of all methods are indicated in boldface.
Method Weibo Reddit Disney Books Enron Avg.
DiffGAD(MLP) 90.0 $\pm$ 1.1 50.6 $\pm$ 0.0 49.1 $\pm$ 0.2 50.0 $\pm$ 0.0 81.9 $\pm$ 5.2 64.3 $\pm$ 1.3
DiffGAD(VAE) 90.5 $\pm$ 0.6 50.6 $\pm$ 0.0 48.5 $\pm$ 0.2 67.1 $\pm$ 2.9 66.0 $\pm$ 1.8 64.5 $\pm$ 1.1
DiffGAD(VGAE) 92.5 $\pm$ 0.0 56.1 $\pm$ 0.0 52.1 $\pm$ 3.2 51.0 $\pm$ 1.1 57.8 $\pm$ 0.2 61.9 $\pm$ 0.5
DiffGAD(GTrans) 93.0 $\pm$ 0.7 56.2 $\pm$ 0.1 51.4 $\pm$ 0.6 51.5 $\pm$ 0.5 65.7 $\pm$ 4.9 63.6 $\pm$ 1.4
DiffGAD(GAE) 93.4 $\pm$ 0.3 56.3 $\pm$ 0.1 54.5 $\pm$ 0.2 66.4 $\pm$ 1.8 71.6 $\pm$ 7.0 68.4 $\pm$ 1.9
In this subsection, we evaluate the adaptation ability of DiffGAD with different backbones, Specifically, we utilize 4 different encoder & decoder architectures, including VGAE, MLP, VAE, Graph Transformer, and evaluate the ROC-AUC result. The experimental results on the backbones are listed in Table 11, and results on DiffGAD with different backbones are shown in Table 12. From the experimental results, we can draw following observations:
  1. 1.
    GAE and DiffGAD(GAE) consistently exhibit superior performance compared to different encoder & decoder architectures across different datasets in terms of ROC-AUC. Such results underscore the robustness of GAE and the generalization ability of DiffGAD.
  2. 2.
    Without encoding graph structure, the DiffGAD(MLP) achieves notable results on the Enron dataset, this indicates that the graph structure can lead to a loss of discriminative information, as MLPs rely solely on node features as input.
  3. 3.
    The different performance of GAE, VAE, and VGAE reflects the impact of different encoder and decoder architectures. However, the architecture of Graph Transformer may be somewhat excessive for this task, as it exhibits a relatively high complexity, but without demonstrating any performance improvement.
To sum up: we find that (1) DiffGAD achieves robust performance gains over different architectures, which demonstrates its adaptation ability, and (2) a simple architecture (i.e. GAE) is enough for latent space construction over current datasets, which preserves sufficient discriminative information, and further works with this latent space can boost the detection performance.
C.6 Choices about $\lambda$.
Table 13: The homophily and heterophily of anomalies on 5 datasets.
Weibo Reddit Disney Books Enron
Homophily 0.87 0.18 0.01 0.00 1.00
Heterophily 0.13 0.82 0.99 1.00 0.00
In this subsection, we describe the selection of the key hyper-parameter $\lambda$. Specifically, we statistic the heterophily of the Abnormal samples over different dataset, and the results are listed in Table 13. We observe that on the dataset with larger anomaly heterophily, such as Enron, Disney, and Books, larger $\lambda$ works well, while on the dataset with smaller anomaly heterophily, like Weibo and Reddit, smaller $\lambda$ works well. To summarize, we draw the following conclusion:
  1. 1.
    Larger $\lambda$ works better on the datasets where anomaly samples has larger heterophily, and we recommend using $\lambda$ larger than 1 on such data. Larger heterophily for anomaly samples means they are hidden within normal samples, and thus they are harder to detect.
  2. 2.
    Smaller $\lambda$ works well on the datasets where anomaly samples has smaller heterophily, and we recommend using $\lambda$ smaller than 1 on such data. Smaller heterophily for anomaly samples means they are clustered together, and thus they are easier to detect.
C.7 Detailed Baseline Description
@@
DOMINANT (Ding et al., 2019): DOMINANT is a pioneering work that leverages GCN and AE for anomaly detection. It follows an encoder-decoder architecture with a two-layer GCN. The node feature decoder is also a two-layer GCN, while the graph structure is decoded by a one-layer GCN and dot product. The anomaly score of each node is determined by the weighted sum of two decoders.